# A dimension reduction approach to edge weight estimation for use in spatial models

### v1

@@ Abstract
Models for areal data are traditionally defined using the neighborhood structure of the regions on which data are observed. The unweighted adjacency matrix of a graph is commonly used to characterize the relationships between locations, resulting in the implicit assumption that all pairs of neighboring regions interact similarly, an assumption which may not be true in practice. It has been shown that more complex spatial relationships between graph nodes may be represented when edge weights are allowed to vary. Christensen and Hoff (2024) introduced a covariance model for data observed on graphs which is more flexible than traditional alternatives, parameterizing covariance as a function of an unknown edge weights matrix. A potential issue with their approach is that each edge weight is treated as a unique parameter, resulting in increasingly challenging parameter estimation as graph size increases. Within this article we propose a framework for estimating edge weight matrices that reduces their effective dimension via a basis function representation of of the edge weights. We show that this method may be used to enhance the performance and flexibility of covariance models parameterized by such matrices in a series of illustrations, simulations and data examples.
@@ 1 Introduction
Studies involving areal, or region indexed data are common in many settings including ecology (e.g. Hanks and Hooten, 2013), economics (e.g. Arbia, 2012), public health (e.g. Jin et al., 2005) and sociology (e.g. Garner and Raudenbush, 1991). Models for spatially indexed data are typically defined such that observations close to one another in space exhibit higher correlation than observations which are far apart. In contrast to spatial models for point-indexed data, which tend to characterize between-location dependence as a function of the geographic distances between observations, models for areal data generally define spatial dependence in terms of neighborhood or adjacency structure (Cressie, 1993).
@@
The spatial structure of areal data may be represented using a graph, with each spatial region represented by a graph node and with edges existing between each pair of nodes corresponding to physically adjacent regions. First introduced by Besag (1974), the most commonly used spatial model for areal data is the conditional autoregressive (CAR) model. Given a graph $G=(V,E)$ having $p$ nodes, we observe the random vector y, where $y_{1},...,y_{p}$ are observations at each node of $G$. A CAR model defines a Markov random field where the conditional distribution of each element of y is determined by the values of y at adjacent nodes:
$y_{i}|\textbf{y}_{-i}\sim N\left(\sum_{j\sim i}\alpha w_{ij}y_{j},\sigma^{2}\right).$ (1)
Equation 1 results in a marginal distribution for y of
$\textbf{y}\sim N_{p}(\textbf{0},\sigma^{2}(\textbf{I}_{p}-\alpha\textbf{W})^{-1})$ (2)
@@
where $j\sim i$ indicates that node $j$ is adjacent to node $i$, W is a $p\times p$ adjacency or weights matrix (where $w_{ij}>0$ if $j\sim i$ else $w_{ij}=0$), $\sigma^{2}>0$ is a variance parameter, and $\alpha$ controls the degree of spatial correlation. In order to produce a valid covariance matrix, $\alpha$ must be bounded between the largest and smallest eigenvalue of W (Ver Hoef et al., 2018). An alternative form of the CAR model may be written
$\textbf{y}\sim N_{p}(\textbf{0},\sigma^{2}(\text{diag}(\textbf{W}\textbf{1}_{p})-\kappa\textbf{W})^{-1})$ (3)
where $\text{diag}(\cdot)$ returns a diagonal matrix with entries equal to the input vector. This model is valid for all $|\kappa|<1$. The intrinsic conditional autoregressive (ICAR) model is a special case of Equation 3 when $\kappa=1$. This results in an improper distribution, but it may be reconfigured as a $p-1$ dimensional distribution using the constraint $\sum y_{i}=0$ (Besag and Kooperberg, 1995).
@@
While many variants of the CAR and other similarly defined spatial autoregressive models have been developed and utilized, Ver Hoef et al. (2018) notes that in practice almost all such models have been implemented using the unweighted adjacency matrix of $G$, with $w_{ij}=1$ if $i\sim j$, and $w_{ij}=0$ otherwise. The CAR model is frequently used to characterize spatial random effects or an “error process” within a larger generalized linear model (Besag et al., 1991); in many such contexts, interest in the covariance structure of the model has been secondary to inference on the fixed effects governing the model’s mean structure. However, the lack of a more complex weighting scheme results in a pattern of spatial dependence where the partial correlations between all pairs of adjacent regions are identical regardless of location within the overall spatial domain. Such models may be thought of as discrete space equivalents to stationary and isotropic covariance models for point indexed data in continuous space, properties which are frequently inappropriate for real world data (Guttorp et al., 1994).
@@
Christensen and Hoff (2024) introduced a covariance model (which we will refer to as the graph deformation model or GDEF model within this article) for areal data exhibiting greater flexibility than the traditional CAR model. Given a graph $G=(V,E)$ with $p$ nodes and random vector y consisting of observations at each node of $G$, they define $\Sigma=\text{Cov}(\textbf{y})$ as follows:
$\Sigma_{ij}$ $=\sigma^{2}\rho_{\nu}(d_{ij})$ (4)
$d_{ij}$ $=\sqrt{(\textbf{e}_{i}-\textbf{e}_{j})^{\top}\{\textbf{L}^{+}\}^{2}(\textbf{e}_{i}-\textbf{e}_{j})}$
L $=\text{diag}(\textbf{W}\textbf{1}_{p})-\textbf{W}$
@@
Here $\sigma^{2}$ is a scale parameter and $\rho_{\nu}(\cdot)$ is the Matérn correlation function (Matérn, 1960) with smoothness parameter $\nu$. The term $d_{ij}$ may be thought of as the latent distance between nodes $i$ and $j$ and is defined as a function of unknown edge weights matrix W. The matrix $\{\textbf{L}^{+}\}^{2}=\textbf{L}^{+}\textbf{L}^{+}$ is the square of the generalized inverse of the graph Laplacian L, and $\{\textbf{e}_{i}\}_{1:p}$ are the standard basis vectors. The edge weights $\{w_{ij}\}_{i\sim j}$ are estimated from the observed data.
@@
The GDEF model described in Equation 4 results in a valid covariance matrix for any combination of positive edge weights and the matrix W is an identifiable parameter. While Christensen and Hoff (2024) demonstrate the flexibility and interpretive value of the GDEF approach, the authors also acknowledge that the model has limited scalability when applied to large graphs given the number of model parameters (each non-zero edge weight must be estimated individually) and computational intensity of the Markov chain Monte Carlo (MCMC) algorithm they used to obtain parameter estimates.
Both CAR and GDEF models are constructed using the edge weights matrix W. While the GDEF model could be improved by coming up with a simpler and more efficient approach to estimation of W, CAR models may be improved in practice by using a more complex approach. As noted, CAR models have traditionally been defined with W equal to the unweighted adjacency matrix of the graphical representation of the spatial domain, a choice that is both simple and convenient but may be restrictive depending on the data and scientific question of interest. Despite likely under-utilization in practice, more complex weighting schemes for CAR models have been shown to be valuable, as in Hanks and Hooten (2013) and Ejigu and Wencheko (2020) which define edge weights using linear combinations of environmental covariates. Weighting schemes with even greater flexibility are also possible and potentially enable deeper understanding of the interaction between discretely observed spatial processes and the the spatial domain over which they are observed.
@@
Within this article we introduce a new approach to estimating edge weights matrices that may be used within the GREF framework of Christensen and Hoff (2024) and which allows for the specification of CAR models that are considerably more flexible than existing variants. In the following section we provide a brief overview of necessary background information related to spatial basis functions, graphs, and spectral theory. In Section 3 we present our method and discuss some of its properties. Section 4 illustrates and evaluates the performance of our approach through a series of simulation studies and a real data example. We conclude the article with a discussion of potential applications and extensions.
@@ 2 Background
Given a simple, undirected graph $G=(V,E)$, random variables $\{y_{v}:v\in V\}$ observed at the nodes $V$ of $G$, and defining $\Sigma_{vv^{\prime}}=\text{Cov}(y_{v},y_{v^{\prime}})$, we parameterize the covariance matrix as a function of W, a matrix of unknown positive edge weights. We define $\mathcal{W}_{G}$ to be the space of all edge weight matrices W possible under $G$ such that $w_{ij}>0$ if an edge exists between nodes $i$ and $j$ and $w_{ij}=0$ otherwise. As the number of nodes in $G$ increases, so too does the size of $\mathcal{W}_{G}$. As such, setting a prior distribution for W over the complete space $\mathcal{W}_{G}$ (as in Christensen and Hoff (2024)) may not be feasible for graphs containing more than a few hundred nodes. To address this, we restrict the parameter space to $\textbf{W}\in\mathcal{W}^{B}_{G}\subset\mathcal{W}_{G}$, where $\mathcal{W}^{B}_{G}$ contains all possible edge weights matrices that can be obtained from a basis function representation (to be described in detail later) of the non-zero edge weights.
@@ 2.1 Basis functions and spatial models
$y(\textbf{s})=\textbf{x}^{\top}(\textbf{s})\beta+u(\textbf{s})+\epsilon(\textbf{s})$ (5)
where $\beta$ is a vector of coefficients, $u(\textbf{s})$ is a spatial random effect and $\epsilon(\textbf{s})\overset{iid}{\sim}N(0,\sigma^{2})$ for all $\textbf{s}\in\mathcal{D}$. Let $\textbf{u}=(u(\textbf{s}_{1}),...,u(\textbf{s}_{p}))$ denote the vector of spatial random effects at some collection of $p$ locations.
Instead of modeling $\textbf{u}\sim N_{p}(\textbf{0},\Sigma)$ it is common to define $\textbf{u}=\textbf{M}\eta$, where M is an $p\times k$ matrix of $k<<p$ basis functions, and $\eta\sim N_{k}(\textbf{0},\Phi)$. Reduced-rank statistical models representing spatial random effects using basis functions are widespread. Examples of this include, but are not limited to the kernel convolution approach (Higdon, 1998), predictive processes (Banerjee et al., 2008), spatial splines (Sangalli et al., 2013) the stochasistic partial differential equation approach (Lindgren et al., 2011), and various eigenvector and principle component based methods (Hughes and Haran, 2013; Higdon et al., 2008; Lee and Haran, 2022). Note that the majority of these methods were developed for point indexed data, though the framework of Hughes and Haran (2013) is designed specifically for areal data. Applying methods designed for one data type to another is possible, but requires assumptions that may not be appropriate in all settings (Gelfand et al., 2001).
Basis function methods tend to be very computationally efficient relative to full-rank Gaussian process models. Notably, some approaches to basis functions specification are also capable of addressing issues such as nonstationarity and anisotropy (Schmidt and Guttorp, 2020). It has often been observed that methods based on low-rank representations of spatial random effects may result in oversmoothing as well as bias in the estimation of variance components (Stein, 2014; Datta et al., 2016; Banerjee et al., 2010). Additionally, when pre-specifying the design matrix M using some set of basis functions, the covariance of u is restricted to matrices of the form $\textbf{M}\Phi\textbf{M}^{\top}$: in many instances, $\Phi$ is assumed to be a diagonal matrix or multiple of the identity matrix, resulting in strong and potentially inappropriate assumptions about model covariance structure. This may be especially true if analysis goals go beyond spatial prediction and smoothing and include inference regarding model covariance. Rather than modelling spatial random effects for areal data as a linear combination of basis functions, we propose using basis functions in a novel parameterization of edge weights matrix W which is then incorporated into a CAR or GDEF model, resulting in a class of full-rank covariance matrices that are more flexible than traditional autoregressive models and less computationally demanding than the implementation of the GDEF used by Christensen and Hoff (2024).
@@ 2.2 Line graphs and an eigenvector basis
We now describe the basis used in this article. Given a graph $G=(V,E)$ having $p$ nodes and $q$ edges, one may define its line graph $L(G)$ as a graph with $q$ nodes corresponding to each of the edges of $G$, and edges between each pair of nodes corresponding to coincident edges in the original graph $G$. (Edges are considered coincident if they share a common node). Figure 1 illustrates the relationship between a graph $G$ and its line graph $L(G)$ for a graph containing five nodes.
@@
Figure 1: The graph $G$ has corresponding line graph $L(G)$
Let $\textbf{A}^{L}$ be the $q\times q$ adjacency matrix of $L(G)$, where $a^{L}_{ij}=1$ if nodes $i$ and $j$ of $L(G)$ are adjacent and $a^{L}_{ij}=0$ otherwise. Let $\textbf{L}^{L}=\text{diag}(\textbf{A}^{L}\textbf{1}_{q})-\textbf{A}^{L}$ be the Laplacian matrix of $L(G)$. If $\textbf{V}\mbox{\boldmath$\Lambda$}\textbf{V}^{\top}$ is the eigendecomposition of $\textbf{L}^{L}$, the columns vectors of V form an orthonormal basis for $\mathbb{R}^{q}$. The eigenvalues of Laplacian matrices are non-negative and generally ordered from smallest to largest, with smallest non-zero eigenvalues corresponding to the most “important” eigenvectors. Laplacian matrices for connected graphs are positive-semidefinite and have exactly one eigenvalue (denoted $\lambda_{1}$) equal to 0 with corresponding eigenvector $\textbf{v}_{1}\propto\textbf{1}$. Despite some referring to $\textbf{v}_{1}$ as the “trivial” eigenvector, it has interpretive value as an intercept column within an eigenvector basis. Considerable work has been done in the field of spectral graph theory investigating the properties of the eigenvalues and eigenvectors of Laplacian matrices. See Spielman (2012) for greater detail on this topic.
To visualize the basis characterized by the eigendecomposition of the line graph Laplacian (LGL) Figure 2 depicts a selection of LGL eigenvectors for a $30\times 30$ lattice graph, while Figure 3 depicts the LGL eigenvectors for the graph of the 100 counties of North Carolina (the spatial domain analyzed in Christensen and Hoff (2024)). Note that each vector element corresponds to a specific edge of the original graph. We also highlight the harmonic behavior of these eigenvectors, with later eigenvectors exhibiting higher frequency oscillations between positively and negatively valued regions of the graph. This makes selection of the first $k$ eigenvectors (corresponding to the $k$ smallest eigenvalues) a natural choice of basis, with the “most important” initial eigenvectors capturing larger-scale trends in edge weight behavior and additional eigenvectors adding finer detail.
Figure 2: A selection of LGL eigenvectors for $30\times 30$ lattice grid
Figure 3: A selection of LGL eigenvectors for graph of the 100 counties of North Carolina

### v2

@@ Abstract
Models for areal data are traditionally defined using the neighborhood structure of the regions on which data are observed. The unweighted adjacency matrix of a graph is commonly used to characterize the relationships between locations, resulting in the implicit assumption that all pairs of neighboring regions interact similarly, an assumption which may not be true in practice. It has been shown that more complex spatial relationships between graph nodes may be represented when edge weights are allowed to vary. Christensen and Hoff, (2024) introduced a covariance model for data observed on graphs which is more flexible than traditional alternatives, parameterizing covariance as a function of an unknown edge weights matrix. One potential issue with their approach is that each edge weight is treated as a unique parameter, resulting in increasingly challenging parameter estimation as graph size increases. Within this article we propose a framework for estimating edge weight matrices, reducing their effective dimension via a basis function representation. By further leveraging fast matrix calculations for derivative expressions, we enhance the performance and flexibility of covariance models parameterized by such matrices, and demonstrate the utility of our method in a series of illustrations, simulations and data examples.
Keywords: Areal data, covariance, line graph, eigenvector basis, CAR model, spatial deformation
@@ 1 Introduction
Studies involving areal, or region indexed data are common in many settings including ecology (e.g. Hanks and Hooten,, 2013), economics (e.g. Arbia,, 2012), public health (e.g. Jin et al.,, 2005) and sociology (e.g. Garner and Raudenbush,, 1991). Models for spatially indexed data are typically defined such that observations close to one another in space exhibit higher correlation than observations which are far apart Stein et al., (2004). In contrast to spatial models for point-indexed data, which tend to characterize between-location dependence as a function of the geographic distances between observations, models for areal data generally define spatial dependence in terms of neighborhood or adjacency structure (Cressie,, 1993).
@@
The spatial structure of areal data may be represented using a graph, with each spatial region as a graph node and with edges existing between each pair of nodes corresponding to physically adjacent regions. First introduced by Besag, (1974), the most commonly used spatial model for areal data is the conditional autoregressive (CAR) model. Given vertex set $V$ and edge set $E$, we define the graph $G=(V,E)$ with $p=|V|$ nodes. We observe the random vector y, where $y_{1},...,y_{p}$ are observations at each node of $G$. A CAR model for y defines a Markov random field where the conditional distribution of each element of y is determined by the values of y at adjacent nodes:
$y_{i}|\textbf{y}_{-i}\sim N\left(\sum_{j\sim i}\alpha w_{ij}y_{j},\sigma^{2}\right),$ (1)
where $j\sim i$ indicates that node $j$ is adjacent to node $i$, $w_{ij}$ characterizes the strength of the relationship between nodes $i$ and $j$, $\sigma^{2}>0$ is a variance parameter, and $\alpha$ controls the degree of spatial correlation. Equation 1 results in a marginal distribution for y of
$\textbf{y}\sim N_{p}(\textbf{0},\sigma^{2}(\textbf{I}_{p}-\alpha\textbf{W})^{-1}),$ (2)
@@
where W is a $p\times p$ adjacency or weights matrix defined such that $w_{ij}>0$ if $j\sim i$ and $w_{ij}=0$ otherwise. In order to produce a valid covariance matrix, $\alpha$ must be bounded between the largest and smallest eigenvalue of W (Ver Hoef et al.,, 2018). An alternative form of the CAR model is
$\textbf{y}\sim N_{p}(\textbf{0},\sigma^{2}(\text{diag}(\textbf{W}\textbf{1}_{p})-\kappa\textbf{W})^{-1}),$ (3)
[…]
@@
While many variants of the CAR model and other similarly defined spatial autoregressive models have been developed and utilized, Ver Hoef et al., (2018) note that in practice almost all such models have been implemented using the unweighted adjacency matrix of $G$, with $w_{ij}=1$ if $i\sim j$, and $w_{ij}=0$ otherwise. The CAR model is frequently used to characterize spatial random effects or an “error process” within a larger generalized linear model (Besag et al.,, 1991); in many such contexts, interest in the covariance structure of the model has been secondary to inference on the fixed effects governing the model’s mean structure. However, the lack of a more complex weighting scheme results in a pattern of spatial dependence where the partial correlations between all pairs of adjacent regions are near-identical regardless of location within the overall spatial domain. Such models may be thought of as discrete space equivalents to stationary and isotropic covariance models for point indexed data in continuous space, properties which are frequently inappropriate for real world data (Guttorp et al.,, 1994). Despite likely under-utilization in practice, more complex weighting schemes for CAR models have been shown to be useful, as in Wall, (2004) who examines the implied correlation structures in various models for areal data, and Hanks and Hooten, (2013) and Ejigu and Wencheko, (2020) who define edge weights using linear combinations of environmental covariates. Weighting schemes with even greater flexibility are also possible and they potentially enable more realistic interaction structures in the discretely observed spatial processes.
@@
Christensen and Hoff, (2024) introduced a covariance model for areal data exhibiting greater flexibility than the traditional CAR model. We refer to this as the graph deformation (GDEF) model. Given the graph $G=(V,E)$, they define $\Sigma=\text{Cov}(\textbf{y})$ as follows:
$\Sigma_{ij}$ $=\sigma^{2}\rho_{\nu}(d_{ij}),$ (4)
$d_{ij}$ $=\sqrt{(\textbf{e}_{i}-\textbf{e}_{j})^{\top}\{\textbf{L}^{+}\}^{2}(\textbf{e}_{i}-\textbf{e}_{j})},$
L $=\text{diag}(\textbf{W}\textbf{1}_{p})-\textbf{W},$
@@
Here $\sigma^{2}$ is a scale parameter and $\rho_{\nu}(\cdot)$ is the Matérn correlation function (Matérn,, 1960) with smoothness parameter $\nu$. The term $d_{ij}$ may be thought of as the latent distance between nodes $i$ and $j$ and is defined as a function of unknown edge weights matrix W. The matrix $\{\textbf{L}^{+}\}^{2}=\textbf{L}^{+}\textbf{L}^{+}$ is the square of the generalized inverse of the graph Laplacian L, and $\textbf{e}_{i}$, $i=1,\ldots,p$ are the standard basis vectors. The edge weights $\{w_{ij}\}_{i\sim j}$ are estimated from the observed data and may be interpreted as the strength of connectivity between adjacent nodes or regions. Related models have been suggested for resistance distances in physics and chemistry (Klein and Randić,, 1993).
@@
The GDEF model described in Equation 4 results in a valid covariance matrix for any combination of positive edge weights and the matrix W is an identifiable parameter. While Christensen and Hoff, (2024) demonstrate the flexibility and interpretive value of the GDEF approach, they acknowledge that the model has limited scalability when applied to large graphs and using their Markov chain Monte Carlo (MCMC) algorithm for estimation.
A key question for both the implementation of the GDEF model and within the literature concerning more complex weighting schemes for spatial autoregressive models is how to define and estimate the edge weights matrix W. The individual estimation of each unique edge weight, as in Christensen and Hoff, (2024), proves to be computationally impractical for graphs with more than a few hundred nodes, while also introducing potential identifiability issues for models that have not been carefully specified. On the other hand, the covariance structures resulting from previously proposed edge weighting schemes, such as defining edge weights as linear combinations of covariates, or scaling edge weights to be proportional to distances between region centroids or the length of the border shared by two regions, may not be flexible enough to accurately characterize some spatial processes.
The principle contributions from our work within this article are as follows:
  • We introduce a dimension reduction approach for the specification and estimation of an unknown edge weights matrix that utilizes eigenvectors from the Laplacian matrix of a network’s line graph as a basis. This method allows for the specification of CAR and other autoregressive models that are considerably more flexible than existing variants, while significantly improving the practical usability of the GDEF framework.
  • By embedding the basis function representation of W within a covariance model rather than directly defining spatial random effects as linear combinations of basis functions as has traditionally been done, we leverage the computational advantages of such approaches while producing a class of models with fewer restrictions on the imposed correlation structure.
  • By incorporating our framework for edge weight estimation within the GDEF framework, we produce a nonstationary spatial covariance model that may be interpreted through the lens of the spatial deformation approach of Sampson and Guttorp, (1992) that may be fit with no replicate observations of the spatial process.
@@
In Section 2 of the article we provide a brief overview of necessary background information related to spatial basis functions, graphs, and spectral theory. In Section 3 we present our approach to edge weight estimation and illustrate some of its properties when incorporated into existing methods. In Section 4 we illustrate and evaluate the performance of our approach through a series of simulation studies and a real data example. We conclude the article with a discussion of potential applications and extensions.
@@ 2 Background
Given a simple, undirected graph $G=(V,E)$, random variables $y_{i}$, $i\in V$, and defining $\Sigma_{ij}=\text{Cov}(y_{i},y_{j})$, we parameterize the covariance matrix $\Sigma$ as a function of W, a matrix of unknown positive edge weights. We define $\mathcal{W}_{G}$ to be the space of all edge weight matrices W that may be defined given $G$ such that $w_{ij}>0$ if an edge exists between nodes $i$ and $j$ and $w_{ij}=0$ otherwise. Note that this means that when we speak of estimating W, we are referring only to the estimation of the edge weights within a fixed network structure, rather than trying to determine the existence or nonexistence of edges between observed nodes. As the number of nodes in $G$ increases, so too does the size of $\mathcal{W}_{G}$. To address this challenge, we here restrict the parameter space to $\textbf{W}\in\mathcal{W}^{B}_{G}\subset\mathcal{W}_{G}$, where $\mathcal{W}^{B}_{G}$ contains all possible edge weights matrices that can be obtained from a basis function representation of the non-zero edge weights.
@@ 2.1 Basis functions and spatial models
$y(\textbf{s})=\textbf{x}^{\top}(\textbf{s})\beta+z(\textbf{s})+\epsilon(\textbf{s}),$ (5)
[…]
Let $\textbf{z}=(z(\textbf{s}_{1}),...,z(\textbf{s}_{p}))$ denote the vector of spatial random effects at some collection of $p$ locations. Instead of modeling $\textbf{z}\sim N_{p}(\textbf{0},\Sigma)$ it is common to define $\textbf{z}=\textbf{M}\eta$, where M is an $p\times k$ matrix of $k\ll p$ basis functions, and $\eta\sim N_{k}(\textbf{0},\Phi)$. Reduced-rank statistical models representing spatial random effects using basis functions are widespread. Examples of this include, but are not limited to the kernel convolution approach (Higdon,, 1998), predictive processes (Banerjee et al.,, 2008), spatial splines (Sangalli et al.,, 2013), the stochastic partial differential equation approach (Lindgren et al.,, 2011), and various eigenvector and principle component based methods (Higdon et al.,, 2008; Hughes and Haran,, 2013; Lee and Haran,, 2022). Note that the majority of these methods were developed for point indexed data, though the framework of Hughes and Haran, (2013) is designed specifically for areal data. Applying methods designed for one data type to another is possible, but requires assumptions that may not be appropriate in all settings (Gelfand et al.,, 2001).
Basis function methods tend to be very computationally efficient relative to full-rank Gaussian process models. Notably, some approaches to basis functions specification are also capable of addressing issues such as nonstationarity and anisotropy (Schmidt and Guttorp,, 2020). It has often been observed that methods based on low-rank representations of spatial random effects may result in oversmoothing as well as bias in the estimation of variance components (Banerjee et al.,, 2010; Stein,, 2014; Datta et al.,, 2016). Additionally, when pre-specifying the design matrix M using some set of basis functions, the covariance of u is restricted to matrices of the form $\textbf{M}\Phi\textbf{M}^{\top}$: in many instances, $\Phi$ is assumed to be a diagonal matrix or multiple of the identity matrix, resulting in strong and potentially inappropriate assumptions about model covariance structure. This may be especially true if analysis goals go beyond spatial prediction and smoothing and include inference regarding model covariance. Rather than modeling spatial random effects for areal data as a linear combination of basis functions, we propose using basis functions in a novel parameterization of edge weights matrix W which is then incorporated into a CAR or GDEF model, resulting in a flexible class of full-rank covariance matrices that are less computationally demanding than the implementation used by Christensen and Hoff, (2024).
@@ 2.2 Line graphs and an eigenvector basis
For a graph $G$, one may define its line graph $L(G)$ as a graph with $q$ nodes corresponding to each of the edges of $G$, and edges between each pair of nodes corresponding to coincident edges in the original graph $G$. (Edges are considered coincident if they share a common node). Figure 1 illustrates the relationship between a graph $G$ and its line graph $L(G)$ for a graph containing $p=5$ nodes and $q=8$ edges.
@@
Figure 1: The graph $G$ has corresponding line graph $L(G)$. The five nodes are $v_{1},\ldots,v_{5}$. The eight edges are $e_{1},\ldots,e_{8}$.
Let $\textbf{A}^{L}$ be the $q\times q$ adjacency matrix of $L(G)$, where $a^{L}_{ij}=1$ if nodes $i$ and $j$ of $L(G)$ are adjacent and $a^{L}_{ij}=0$ otherwise. Let $\textbf{L}^{L}=\text{diag}(\textbf{A}^{L}\textbf{1}_{q})-\textbf{A}^{L}$ be the Laplacian matrix of $L(G)$. If $\textbf{V}\mbox{\boldmath$\Lambda$}\textbf{V}^{\top}$ is the eigendecomposition of $\textbf{L}^{L}$, the columns vectors of V form an orthonormal basis for $\mathbb{R}^{q}$. The eigenvalues of Laplacian matrices are non-negative and generally ordered from smallest to largest, with smallest non-zero eigenvalues corresponding to the most “important” eigenvectors. Laplacian matrices for connected graphs are positive-semidefinite and have exactly one eigenvalue (denoted $\lambda_{1}$) equal to 0 with corresponding eigenvector $\textbf{v}_{1}\propto\textbf{1}$. Despite some referring to $\textbf{v}_{1}$ as the “trivial” eigenvector, it has interpretive value as an intercept column within an eigenvector basis. Considerable work has been done in the field of spectral graph theory investigating the properties of the eigenvalues and eigenvectors of Laplacian matrices (Spielman,, 2012).
Laplacian matrices have many applications in various types of data analysis. Within the field of signal processing for data on undirected graphs, the matrix $\textbf{V}^{-1}$ defines the graph Fourier transform, and is used to decompose signals in terms of the eigenvectors of the Laplacian (Sandryhaila and Moura,, 2013; Ortega et al.,, 2018). In the field of spatial statistics, the eigenvectors of graph Laplacians have been used as a basis to define a smoothly varying processes over networks, as in Upton and Carvalho, (2024) who model urban crime on a street network by estimating node-level regression coefficients in such a manner. One of the most novel contributions of our work in this article is the utilization of the eigenvectors of the Laplacian matrix of a network’s line graph to define a process over its edges rather than its nodes. Similar methods were applied recently in Josephs and Upton, (2025), who estimated the efficacy of different NBA lineups by utilizing a hypergraph representation of all players and player groupings, though the nature of the networks they analyzed and the random variables modeled differ considerably from those in focus here.
In order to understand how the eigendecomposition of the line graph Laplacian (LGL) provides a useful basis for modeling the edge weights within a network, we visualize the the original graph with edges weighted by the the corresponding elements of the LGL eigenvectors. Figure 2 depicts a selection of LGL eigenvectors for a $\times 30$ 30 lattice graph, while Figure 3 shows that of the 100 counties of North Carolina. Note the harmonic behavior of these eigenvectors, with later eigenvectors exhibiting higher frequency oscillations between positively and negatively valued regions of the graph. This makes selection of the first $k$ eigenvectors (corresponding to the $k$ smallest eigenvalues) a natural choice of basis, with the “most important” initial eigenvectors capturing larger-scale trends in edge weight behavior.
Figure 2: A selection of LGL eigenvectors for the $\times 30$ 30 lattice grid.
Figure 3: A selection of LGL eigenvectors for the graph of the 100 counties of North Carolina.